ADNAN MUHAMMED

Grade 12,

Two blocks of mass 2kg and 4 ku are accelerated with same acceleration by a force 10N on a smooth horizontal surface. Then the spring force between the two blocks will be ( spring is massless)

Two blocks of mass 2kg and 4 ku are accelerated with same acceleration by a force 10N on a smooth horizontal surface. Then the spring force between the two blocks will be ( spring is massless)

Grade:12th pass

1 Answers

Arun
6 years ago
the force F = (2 + 4) * a
a = F/ 6
a = 10/6
let the force of spring be T, so
T = 2a
T = 2 * 10/6 = 10/3
hence spring force will be 10/3 Newton